OpenStudy (anonymous):

Solve by the substitution method 4x+6y=-14 -2x+y=19

OpenStudy (anonymous):

Solve the second equation for y. Plug what you get for y into the first equation.

OpenStudy (anonymous):

i have an idea. multiply the second equation by two. leave the first one alone. you get \[4x+6y=-14\] \[-4x+2y=38\] then add to get \[8y=24\] so \[y=3\]

OpenStudy (anonymous):

That's not the substitution method unfortunately. That's the elimination method. Silly teachers imposing silly restrictions on how to do it.

OpenStudy (anonymous):

once you know \[y=3\] replace y by 3 in either equation to find x. i choose the second one \[-2x+3=-19\] \[-2x=-22\] \[x=11\]

OpenStudy (anonymous):

oooooooooooooooooh sorry

OpenStudy (anonymous):

i should really learn to read better. you are right but that method is rather silly here isn't it?

OpenStudy (anonymous):

I find it silly just about anywhere ;p. I nearly always use the elimination method. Particularly since it transitions so well into matrices.

OpenStudy (anonymous):

Can you tell me if the solution, as an ordered pair is (8,3)